Equation of cosine function with period of 3 pi
Answer by lwsshak3(11628)

Equation of cosine function with period of 3 pi
Equation of cos function: y=Acos(Bx-C)+D, A=amplitude, period=2π/B, phase shift=C/B, D=vertical shift.
For period of 3π:
3π=2π/B
B=2/3
Equation of cos function with period of 3 π: y=Acos(2x/3-C)+D